OpenStudy (anonymous):

how do i find (fog)(5) when f(x)=-4x+11, g(x)=2x-3

OpenStudy (anonymous):

Try this hint...I have placed my response on another site since it is easier to type math over there. http://www.demonstranda.com/book/questionDetails/326

OpenStudy (amistre64):

fog should be read as: f, composed of, g; the (5) refers to the value to give to x f(g) = -4g+11 ; and g(x) = 2x-3; at x = 5

OpenStudy (agent0smith):

fog(5) means f( g(5) ) this means you need to find the value of g(5) then plug THAT number into f(x)

OpenStudy (anonymous):

How would i find the value of g(5)?

OpenStudy (agent0smith):

g(x)=2x-3 so plug in 5 in place of x, into 2x-3

OpenStudy (anonymous):

i like to think of the functional notation like this: \[f(x) = -4(x)+11\] that way, if anything replaces x in the parentheses on the left, it does that same in the parentheses on the right. so \[f\circ g(x) = f \left( g(x) \right)= -4\left( g(x) \right)+11=-4(2x-3)+11= -8x+12+11= -8x+23\]

OpenStudy (anonymous):

and then that x value is plugged into the f(x) equation?